OpenStudy (anonymous):

Water is drained out of tank, shaped as an inverted right circular cone that has a radius of 6 cm and a height of 12 cm, at the rate of 3 cm^3/min. At what rate is the depth of the water changing at the instant when the water in the tank is 9 cm deep?

OpenStudy (anonymous):

@CausticSyndicalist @emily0824 @Loser66 @satellite73 @ShadowLegendX @TheSmartOne

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

@satellite73 @amistre64 @aaronq @Callisto @Compassionate @Zarkon @TheSmartOne @sammixboo @pooja195 @nincompoop @sleepyjess @bohotness @Data_LG2 @dtan5457 @Somy

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

What is the volume of water at a given depth "h"? ConeVolume = V(h,r) = \(\dfrac{1}{3}\pi r^{2}h\) Can you eliminate the 'r' by using the triangular, cross-sectional area of the upside down cone? We need V(h).

OpenStudy (anonymous):

Water is drained out of tank, shaped as an inverted right circular cone that has a radius of 6 cm and a height of 12 cm, at the rate of 3 cm^3/min. At what rate is the depth of the water changing at the instant when the water in the tank is 9 cm deep?

OpenStudy (tkhunny):

Why are we repeating the question. Answer my question, please. We need V(h), instead of V(h,r). Can you do it? It is a geometry problem.

OpenStudy (anonymous):

@tkhunny i fugured it out

OpenStudy (tkhunny):

What did you do? Share!